Probability
CBSE Class 10 Mathematics · every board question from this chapter, 2023–2026
- For an event E, P(E)+P(Ē)=x, then the value of x^2-3 is:2026
- Two dice are rolled together. The probability that the sum of the numbers obtained is divisible by 6, is:2026
- Three coins are tossed together. The probability of getting exactly one head, is:2026
- Which of the following can not be the probability of an event?2026
- Two different dice are rolled together. The probability that both the obtained numbers are less than 4, is2026
- A card is drawn at random from a well shuffled deck of 52 playing cards. The probability that it is either a ten or a king is2026
- Meena calculates that the probability of her winning the first prize in a lottery is 0.08. If total 800 tickets were sold, the…2026
- Two dice are rolled together. The probability that sum of the numbers obtained is atmost 12, is2026
- A die is thrown once. Probability of getting a number other than 3 is:2026
- Two dice are rolled together. The probability of getting an outcome ( x, y ) where x>y, is2026
- Assertion: If probability of happening of an event is 0.2 p, p > 0, then p can't be more than 5. Reason (R): P(Ē)=1-P(E) for an…2026
- Assertion: The probability that a leap year has 53 Mondays is 2/7. Reason (R): The probability that a non-leap year has 53…2026
- A bag contains 25 balls. Some of them are yellow and others are green. One ball is drawn at random. If probability of getting a…2026
- Two different coins are tossed simultaneously. What is the probability of getting: (i) at least one head? (ii) at most one tail?…2026
- Two different dice are thrown together. Find the probability that the numbers obtained have: (i) even sum, (ii) even product.2026
- A bag contains 30 balls out of which 'm' number of balls are blue in colour. (i) Find the probability that a ball drawn at random…2026
- Two dice are thrown at the same time. Determine the probability that the (i) sum of the numbers on the two dice is 5, and (ii)…2026
- Two dice of different colours are thrown at the same time. Write down all the possible outcomes. What is the probability that:…2026
- A group of friends wanted to play cards with two identical packs together. While shuffling the cards, three cards are dropped.…2026
- If for any event E, P(E)+P(Ē)=q, then the value of q^2-3 is:2025
- The probability of drawing an even prime number out of numbers from 1 to 30 is:2025
- A pair of dice is thrown once. The probability that sum of numbers appearing on top faces is at least 4 is:2025
- The number of red balls in a bag is 10 more than the number of black balls. If the probability of drawing a red ball at random…2025
- If in a lottery, there are 10 prizes and 30 blanks, then the probability of winning a prize is:2025
- A bag contains red balls and black balls in the ratio 3: 7. A ball is drawn at random. The probability that ball so drawn is…2025
- Two coins are tossed simultaneously. The probability of getting atleast one head is2025
- A piggy bank contains ₹ 1 coins and ₹ 2 coins in the ratio 9: 11 respectively. The piggy bank is accidently dropped and a coin…2025
- Letters A to F are mentioned on six faces of a die such that each face has a different letter. Two such dice are thrown…2025
- In a cricket match, a batsman hits the boundary 7 times out of the 42 balls he plays. The probability of his not hitting a…2025
- A card is selected at random from a deck of 52 playing cards. The probability of it being a red face card is:2025
- If all the red face cards are removed from the deck of 52 playing cards, then the probability of getting a black jack from the…2025
- A die is thrown once. The probability of getting a number which is not a factor of 36, is:2025
- A bag contains red coloured, blue coloured and green coloured balls in the ratio 2: 3: 4. A ball is drawn at random from the…2025
- A pair of dice is thrown. The probability that sum of numbers appearing on top faces is at most 10 is:2025
- The probability of getting a composite number greater than 3 on throwing a die is2025
- A card is drawn at random from a pack of 52 cards. What is the probability that the card drawn is a spade or a king?2025
- Assertion: The probability of selecting a number at random from the numbers 1 to 20 is 1. Reason (R): For any event E, if P(E)=1,…2025
- In an experiment of throwing a die, Assertion: Event E_1: getting a number less than 3 and Event E_2: getting a number greater…2025
- The probability of guessing the correct answer of a certain test question is x/12. If the probability of not guessing the correct…2025
- A bag contains cards which are numbered from 5 to 100 such that each card bears a different number. A card is drawn at random.…2025
- All the face cards are removed from the pack of 52 cards and a card is drawn at random from the remaining cards. Find the…2025
- Saima and Aryaa were born in the month of June in the year 2012. Find the probability that: (i) they have different dates of…2025
- A bag contains balls numbered 2 to 91 such that each ball bears a different number. A ball is drawn at random from the bag. Find…2025
- From a pack of 52 cards, all aces and all kings are removed. A card is drawn at random from the remaining cards. Find the…2025
- Renu and Simran were born in the year 2000 which is a leap year. Find the probability that: (i) both have same birthday. (ii)…2025
- While shuffling a pack of 52 cards, one card was accidently dropped. Find the probability that the dropped card (i) is not a face…2025
- Two friends Anil and Ashraf were born in the December month in the year 2010. Find the probability that: (i) they share same date…2025
- The number of red balls in a bag is three more than the number of black balls. If the probability of drawing a red ball at random…2025
- All face cards of spades are removed from a pack of 52 playing cards and the remaining pack is shuffled well. A card is then…2025
- If 65\% of the population has black eyes, 25\% have brown eyes and the remaining have blue eyes, what is the probability that a…2025
- Two dice are rolled together. Find the probability of getting: (i) a multiple of 2 on one and a multiple of 3 on the other die.…2025
- Three unbiased coins are tossed simultaneously. Find the probability of getting: exactly two tails at least one head at most two…2025
- Two dice are thrown at the same time. Determine the probability that the difference of the numbers on the two dice is 2.2025
- Rahul is a lucky charm for his cricket team. He has a jar of cards with numbers from 10 to 74. Before each match, he draws a card…2025
- A bag contains 3 red balls, 5 white balls and 7 black balls. The probability that a ball drawn from the bag at random will be…2024
- The probability of getting a bad egg in a lot of 400 eggs is 0.045. The number of good eggs in the lot is:2024
- If the probability of a player winning a game is 0.79, then the probability of his losing the same game is:2024
- Two dice are thrown together. The probability that they show different numbers is:2024
- The probability of guessing the correct answer to a certain test question is x/6. If the probability of not guessing the correct…2024
- One ticket is drawn at random from a bag containing tickets numbered 1 to 40. The probability that the selected ticket has a…2024
- One card is drawn at random from a well shuffled deck of 52 playing cards. The probability that it is a red ace card, is:2024
- Which of the following is not probability of an event?2024
- Two dice are tossed simultaneously. The probability of getting odd numbers on both the dice is:2024
- Two dice are rolled together. The probability of getting the sum of the two numbers to be more than 10, is2024
- From the data 1, 4, 7, 9, 16, 21, 25, if all the even numbers are removed, then the probability of getting at random a prime…2024
- Two coins are tossed simultaneously. The probability of getting at most one tail is:2024
- What is the probability that a number selected randomly from the numbers 1, 2, 3,..., 15 is a multiple of 4?2024
- A box contains cards numbered 6 to 55. A card is drawn at random from the box. The probability that the drawn card has a number…2024
- If a digit is chosen at random from the digits 1, 2, 3, 4, 5, 6, 7, 8, 9; then the probability that this digit is an odd prime…2024
- From the letters of the word "MOBILE", a letter is selected at random. The probability that the selected letter is a vowel, is:2024
- The probability of throwing a number greater than 2 with a fair die is:2024
- Two dice are rolled together. The probability of getting sum of numbers on the two dice as 2, 3 or 5, is:2024
- For an event E, if P(E)+P(Ē)=q, then the value of q^2-4 is:2024
- Assertion: In a cricket match, a batsman hits a boundary 9 times out of 45 balls he plays. The probability that in a given ball,…2024
- The king, queen and ace of clubs and diamonds are removed from a deck of 52 playing cards and the remaining cards are shuffled. A…2024
- One card is drawn at random from a well shuffled deck of 52 cards. Find the probability that the card drawn (i) is queen of…2024
- In a pack of 52 playing cards one card is lost. From the remaining cards, a card is drawn at random. Find the probability that…2024
- A carton consists of 60 shirts of which 48 are good, 8 have major defects and 4 have minor defects. Nigam, a trader, will accept…2024
- Three coins are tossed simultaneously. What is the probability of getting (i) at least one head? (ii) exactly two tails? (iii) at…2024
- Three unbiased coins are tossed simultaneously. Find the probability of getting: (i) at least one head. (ii) exactly one tail.…2024
- A jar contains 54 marbles, each of which is blue, green or white. The probability of selecting a blue marble at random from the…2024
- BINGO is game of chance. The host has 75 balls numbered 1 through 75. Each player has a BINGO card with some numbers written on…2024
- In a survey on holidays, 120 people were asked to state which type of transport they used on their last holiday. The following…2024
- In a lottery, there are 5 prizes and 20 blanks. The probability of getting a prize is:2023
- Two dice are rolled together. What is the probability of getting a sum greater than 10?2023
- One card is drawn at random from a well shuffled pack of 52 playing cards. The probability that the drawn card is a queen, is:2023
- A bag contains 100 cards numbered 1 to 100. A card is drawn at random from the bag. What is the probability that the number on…2023
- If three coins are tossed simultaneously, what is the probability of getting at most one tail?2023
- A bag contains 5 pink, 8 blue and 7 yellow balls. One ball is drawn at random from the bag. What is the probability of getting…2023
- Two dice are thrown together. The probability of getting the difference of numbers on their upper faces equals to 3 is:2023
- In a single throw of two dice, the probability of getting 12 as a product of two numbers obtained is:2023
- A bag contains 5 red balls and n green balls. If the probability of drawing a green ball is three times that of a red ball, then…2023
- A card is drawn at random from a well shuffled deck of 52 playing cards. The probability of getting a face card is2023
- Two coins are tossed together. The probability of getting at least one tail is:2023
- A box contains 90 discs, numbered from 1 to 90. If one disc is drawn at random from the box, the probability that it bears a…2023
- Probability of happening of an event is denoted by p and probability of non-happening of the event is denoted by q. Relation…2023
- In a survey, it is found that every fifth person has a vehicle. The probability of a person NOT having a vehicle, is2023
- A girl calculates that the probability of her winning the first prize in a lottery is 0.08. If 6000 tickets are sold, how many…2023
- Which of the following numbers cannot be the probability of happening of an event?2023
- Cards bearing numbers 3 to 20 are placed in a bag and mixed thoroughly. A card is taken out of the bag at random. What is the…2023
- A card is drawn at random from a well-shuffled pack of 52 cards. The probability that the card drawn is not an ace is:2023
- In a group of 20 people, 5 can't swim. If one person is selected at random, then the probability that he/she can swim, is2023
- Assertion: The probability that a leap year has 53 Sundays is 2/7. Reason (R): The probability that a non-leap year has 53…2023
- If a fair coin is tossed twice, find the probability of getting 'atmost one head'.2023
- A bag contains 4 red, 3 blue and 2 yellow balls. One ball is drawn at random from the bag. Find the probability that drawn ball…2023
- A middle school decided to run the following spinner game as a fund-raiser on Christmas Carnival. Making Purple: Spin each…2023
- Computer-based learning (CBL) refers to any teaching methodology that makes use of computers for information transmission. At an…2023
- "Eight Ball" is a game played on a pool table with 15 balls numbered 1 to 15 and a "cue ball" that is solid and white. Of the 15…2023
Answers come from CBSE’s own marking schemes. How we source them.